# Data Selection via Optimal Control for Language Models

### v1

@@ G.2 Test Loss Extrapolation with the Scaling Law
Table 7: Scaling law constants by fitting the test losses on the DCLM corpus.
We extrapolate the test losses on the DCLM corpus [48] of the conventionally trained and PDS-trained LMs with the Scaling Law [40, 44]. Following Hoffmann et al. [40], we consider the scaling law with the following form:
@@
where $N$ is the model parameters, $D$ is the trained tokens, and $A$, $B$, $E$, $\alpha$, $\beta$ are constants. We obtain these constants by minimizing the Huber loss [41]:
@@
Hoffmann et al. [40] optimizes Eq. (54) using the LBFGS algorithm [52]. However, we find this algorithm sensitive to the initialization of the parameters to be optimized. Therefore, we apply a two-stage optimization. Specifically, we first fit the following data scaling curves for $N=$160M, 470M, 1B, and $1.7B$ with non-linear least squares from scipy.optimize.curve_fit [^7] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.curve_fit.html, which is much more robust to the initialization:
@@
In Section 3.3 (Data-Constrained Setting), to compute the expected token demand of conventional pre-training to match the performance of PDS ($r=0.25$, 4 Eps.), we solve for $D$ using the constants in Table 3 and use $\frac{D}{4}$ as the token demand, indicating the LM can be trained for 4 epochs as suggested by Muennighoff et al. [59].
G.3 Ablation Studies
@@
The desired data plays an important role in determining the quality of the selected data. We test different downstream datasets to compute $J({\bm{\theta}})$ in PDS and report the model performance in Table 8. The comparison between the results of using LAMBADA and LIMA shows the importance of the downstream data diversity. Instances in LAMBADA mostly come from stories, while LIMA is composed of instruction-response pairs that cover various tasks, which yields better overall performance. When comparing LIMA, OpenWebText [29], and CC, we conclude that data quality is another major concern. Although OpenWebText has been shown to have better scaling factors [8] and used as the target set [12], replacing it with higher quality LIMA further improves performance. Compared to diversity and quality, large sizes of downstream datasets seem less important, because LIMA performs the best with the least instance number.

### v2

@@ I.4 Test Loss Extrapolation with the Scaling Law
Figure 9: Scaling law constants by fitting the test losses on the DCLM corpus.
We extrapolate the test losses on the DCLM corpus (Li et al., 2024) of the conventionally trained and PDS-trained LMs with the Scaling Law (Hoffmann et al., 2022; Kaplan et al., 2020). Following Hoffmann et al. (2022), we consider the scaling law with the following form:
@@
where $N$ is the model parameters, $D$ is the trained tokens, and $A$, $B$, $E$, $\alpha$, $\beta$ are constants. We obtain these constants by minimizing the Huber loss (Huber, 1992):
@@
Hoffmann et al. (2022) optimizes Eq. (58) using the LBFGS algorithm (Liu & Nocedal, 1989). However, we find this algorithm sensitive to the initialization of the parameters to be optimized. Therefore, we apply a two-stage optimization. Specifically, we first fit the following data scaling curves for $N=$160M, 470M, 1B, and $1.7B$ with non-linear least squares from scipy.optimize.curve_fit [^7] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.curve_fit.html, which is much more robust to the initialization:
@@
In Section 3.3 (Data-Constrained Setting), to compute the expected token demand of conventional pre-training to match the performance of PDS ($r=0.25$, 4 Eps.), we solve for $D$ using the constants in Table 3 and use $\frac{D}{4}$ as the token demand, indicating the LM can be trained for 4 epochs as suggested by Muennighoff et al. (2023).
$N=$160M $N=$470M $N=$1B $N=$1.7B $D=50\times 10^{9}$
Conventional 0.990 0.998 0.993 0.996 0.999
PDS 0.992 0.995 0.997 0.993 0.998
Table 9: Correlations ($R^{2}$) of fitting the scaling curves.
Goodness of Fit.
We evaluate the goodness of fit of the scaling curves with respect to the training token size $D$ and model size $N$ respectively, by computing the correlation coefficient $R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i}(y_{i}-\overline{y})^{2}}$, where $y_{i}$ is the ground truth value and $\hat{y}_{i}$ is the prediction.
Regarding the training token size, to make the original problem a linear regression problem, we convert Eq. (57) into
$\log\left(L(N,D)-E-\frac{A}{N^{\alpha}}\right)=\log B-\beta\log D.$ (62)
Then, we consider $\log\left(l_{i}-E-\frac{A}{N^{\alpha}}\right)$ as the ground truth value for regression, where $l_{i}$ is the observed loss, and $\log B-\beta\log D_{i}$ as the prediction. For each $N\in[160M,470M,1B,1.7B]$ we compute an $R^{2}$ respectively. Similarly, regarding the model size, we convert Eq. (57) into
$\log\left(L(N,D)-E-\frac{B}{D^{\beta}}\right)=\log A-\alpha\log N,$ (63)
to compute the corresponding $R^{2}$ that measures its goodness of fit. For simplicity, we only consider the models at the end of training, where $D=50\times 10^{9}$. The results in Table 9 show that the correlations are sufficiently high, suggesting the scaling curve fits the impact from both the data and model sizes very well.
I.5 Ablation Studies
@@
The desired data plays an important role in determining the quality of the selected data. We test different downstream datasets to compute $J({\bm{\theta}})$ in PDS and report the model performance in Table 10. The comparison between the results of using LAMBADA and LIMA shows the importance of the downstream data diversity. Instances in LAMBADA mostly come from stories, while LIMA is composed of instruction-response pairs that cover various tasks, which yields better overall performance. When comparing LIMA, OpenWebText Gokaslan et al. (2019), and CC, we conclude that data quality is another major concern. Although OpenWebText has been shown to have better scaling factors (Bi et al., 2024) and used as the target set (Brown et al., 2020), replacing it with higher quality LIMA further improves performance. Compared to diversity and quality, large sizes of downstream datasets seem less important, because LIMA performs the best with the least instance number.